# Notions of rank and independence in countably categorical theories

### v1

@@ 5. The exchange property and symmetry
    For every finite $B^{\prime}\subseteq M^{\mathrm{eq}}$, if $\mathrm{rk}^{n}(B^{\prime}/C)<\mathrm{rk}^{n}(B/C)$ and $\mathrm{rk}^{n}(A/B^{\prime}C)<\mathrm{rk}^{n}(A/C)$, then $\mathrm{rk}^{n}(B^{\prime}/AC)<\mathrm{rk}^{n}(B^{\prime}/C)$.
Suppose that $\mathrm{rk}^{n}(A/BC)<\mathrm{rk}^{n}(A/C)$. Also let $\beta=\mathrm{rk}^{n}(B/C)$. (We will show that $\mathrm{rk}^{n}(B/AC)<\beta$.) Then there is an $n$-$c$s $b_{1},\ldots,b_{\beta}$ for $B/C$ with core sequence of indices $0=k_{0}<k_{1}<\ldots<k_{\delta}=\beta$. In particuar, $b_{1},\ldots,b_{\beta}\in\mathrm{acl}^{n}(B)$. We divide the argument into two main cases.
@@
Let $t$ be minimal such that the above holds. Towards a contradiction, suppose that $\mathrm{rk}^{n}(B/AC)=\mathrm{rk}^{n}(B/C)=\beta$. Then there is an $n$-$c$s $b^{\prime}_{1},\ldots,b^{\prime}_{\beta}$ for $B/AC$ with core sequence of indices $0=l_{0}<l_{1}<\ldots<l_{\gamma}=\beta$. By Lemma 5.12, $b_{1},\ldots,b_{\beta}$ is also an $n$-$c$s for $B/AC$ with core sequence of indices $0=k_{0}<k_{1}<\ldots<k_{\gamma}=\beta$. By Lemma 5.4 we have $\gamma=\delta$ and $l_{i}=k_{i}$ for all $i=0,\ldots,\delta$. From Lemma 5.4 it follows that if, for any $m\in\{0,\ldots,\delta-1\}$, we define
@@
$b_{\beta}\in\mathrm{acl}^{n}(\{b_{1},\ldots,b_{\beta-1},a_{1},\ldots,a_{s-1}\}\cup C)$
@@
In Section 4 we saw that, even without Assumption 5.1, $\smile$$|^{n}$ has all the properties of an independence relation except, possibly, for the symmetry property. Now we have:
@@ 6. Connection to rosiness
    A formula $\varphi(\bar{x},\bar{a})$ (with all parameters listed by $\bar{a})$) strongly divides over $C$ if $\mathrm{tp}(\bar{a}/C)$ is nonalgebraic (i.e. has infinitely many realizations) and the set of formulas $\{\varphi(\bar{x},\bar{a}^{\prime}):\mathrm{tp}(\bar{a}^{\prime}/C)=\mathrm{tp}(\bar{a}/C)\}$ is $k$-inconsistent (meaning that any set of $k$ of these formulas is inconsistent, in a model of $T^{\mathrm{eq}}$).
@@
Proof. The assumption that $\varphi(x,b,\bar{c})$ strongly divides over $C$ means that $\mathrm{tp}(b/C)$ is nonalgebraic and, for some $k<\omega$, $\{\varphi(x,b^{\prime},\bar{c}):\mathrm{tp}(b^{\prime}/C)=\mathrm{tp}(b/C)\}$ is $k$-inconsistent. As explained in [21, Remark 2.1.2], it follows from a compactness argument that there is $\theta(y,\bar{d})\in\mathrm{tp}(b/C)$ such that
@@
We say that $\smile$$|^{n}$ is symmetric restricted to $M_{n}$ if for all $a,b\in M_{n}$, and $C\subseteq M^{\mathrm{eq}}$, if $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|$}\;\>}b$ then $b\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}a$.
@@
[…]
Recall the conclusion above that $\bar{b}\in\mathrm{acl}^{\mathrm{eq}}(\bar{a}^{\prime}D)\setminus\mathrm{acl}^{\mathrm{eq}}(D)$. Since $\bar{b}\in M_{n}$ it follows that $\bar{b}\in\mathrm{acl}^{n}(\bar{a}^{\prime}D)\setminus\mathrm{acl}^{n}(D)$. Since Assumption 6.5 and Proposition 5.14 imply that $\smile$$|^{n}$ has the symmetry property restricted to $M_{n}$ it follows from Lemma 6.4 that $\bar{a}^{\prime}\underset{D}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}\bar{b}$ which contradicts the conclusion above that $\bar{a}^{\prime}\underset{D}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}$. Hence we conclude that $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}\bar{b}$. $\square$
Corollary 6.7.
Suppose that Assumption 6.5 holds for all $n<\omega$. If $a,b\in M^{\mathrm{eq}}$, $C\subseteq M^{\mathrm{eq}}$ and $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$, then $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$ for all sufficiently large $n<\omega$.
Proof. Let $a,b\in M^{\mathrm{eq}}$ and $C\subseteq M^{\mathrm{eq}}$. There is $n<\omega$ such that $a,b\in M_{n}$. Since $M_{n}\subseteq M_{m}$ whenever $n<m$ the corollary now follows from Lemma 6.6. $\square$
@@
Proof. By Remark 2.8 it suffices to prove that there do not exist $a\in M^{\mathrm{eq}}$ and $b_{\alpha}\in M^{\mathrm{eq}}$, for $\alpha<\aleph_{1}$, such that $a\underset{(b_{i})_{i<\alpha}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\alpha}$ for all $\alpha<\aleph_{1}$. Towards a contradiction, suppose that there are $a\in M^{\mathrm{eq}}$ and $b_{\alpha}\in M^{\mathrm{eq}}$, for $\alpha<\aleph_{1}$, such that $a\underset{(b_{i})_{i<\alpha}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\alpha}$ for all $\alpha<\aleph_{1}$. Let $B=\{b_{\alpha}:\alpha<\aleph_{1}\}$. By Lemma 3.12, there is a countable $C\subseteq B$ such that $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ for all $n<\omega$. Let $\beta=\sup\{\alpha:b_{\alpha}\in C\}$. As $\aleph_{1}$ is a regular cardinal and $C$ is countable it follows that $\beta<\aleph_{1}$, so $\beta$ is a countable ordinal. Let $B_{\beta}=\{b_{\alpha}:\alpha\leq\beta\}$ so $B_{\beta}$ is countable and $C\subseteq B_{\beta}$. By the monotonicity of $\smile$$|^{n}$ we get $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ for all $n<\omega$. By assumption we have $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\beta+1}$, so Corollary 6.7 (which uses Assumption 6.5) implies that $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\beta+1}$ for all sufficiently large $n$. Since $b_{\beta+1}\in B$ it follows that $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}B$ for all sufficiently large $n$. But this contradicts the earlier conclusion that $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ for all $n<\omega$. $\square$
@@
For definitions of superrosy theory and $U^{\text{\rm\th }}$-rank see [7, 21].
@@
If $T$ is $\omega$-categorical with soft elimination of imaginaries and Assumption 6.5 holds for $n=0$, then $T$ is superrosy with finite $U^{\text{\rm\th }}$-rank.
Proof. Let $T$ be $\omega$-categorical with soft elimination of imaginaries and suppose that Assumption 6.5 holds for $n=0$. It follows from Theorem 5.15 that $\smile$$|^{0}$ is an independence relation. Since $T$ has soft elimination of imaginaries Lemma 4.9 implies that, for all $n<\omega$ and all $A,B,C\subseteq M^{\mathrm{eq}}$, $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}B$. Hence, for all $n<\omega$, $\smile$$|^{n}$ is an independence relation. It follows that we can argue just as in the proofs of Lemma 6.6 and its corollary and conclude that if $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$ then $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$ for all sufficiently large $n$.

### v2

@@ 5. The exchange property and symmetry
    For every finite $B^{\prime}\subseteq M_{n}$, if $\mathrm{rk}^{n}(B^{\prime}/C)<\mathrm{rk}^{n}(B/C)$ and $\mathrm{rk}^{n}(A/B^{\prime}C)<\mathrm{rk}^{n}(A/C)$, then $\mathrm{rk}^{n}(B^{\prime}/AC)<\mathrm{rk}^{n}(B^{\prime}/C)$.
Suppose that $\mathrm{rk}^{n}(A/BC)<\mathrm{rk}^{n}(A/C)$. Also let $\beta=\mathrm{rk}^{n}(B/C)$. (We will show that $\mathrm{rk}^{n}(B/AC)<\beta$.) Then there is an $n$-$c$s $b_{1},\ldots,b_{\beta}$ for $B/C$ with core sequence of indices $0=k_{0}<k_{1}<\ldots<k_{\delta}=\beta$. In particular, $b_{1},\ldots,b_{\beta}\in\mathrm{acl}^{n}(B)$. We divide the argument into two main cases.
@@
Let $t$ be minimal such that the above holds. Towards a contradiction, suppose that $\mathrm{rk}^{n}(B/AC)=\mathrm{rk}^{n}(B/C)=\beta$. Then there is an $n$-$c$s $b^{\prime}_{1},\ldots,b^{\prime}_{\beta}$ for $B/AC$ with core sequence of indices $0=l_{0}<l_{1}<\ldots<l_{\gamma}=\beta$. By Lemma 5.12, $b_{1},\ldots,b_{\beta}$ is also an $n$-$c$s for $B/C$ with core sequence of indices $0=k_{0}<k_{1}<\ldots<k_{\gamma}=\beta$. By Lemma 5.4 we have $\gamma=\delta$ and $l_{i}=k_{i}$ for all $i=0,\ldots,\delta$. From Lemma 5.4 it follows that if, for any $m\in\{0,\ldots,\delta-1\}$, we define
@@
$b_{\beta}\in\mathrm{acl}^{n}(\{b_{1},\ldots,b_{\beta-1},a_{1},\ldots,a_{s}\}\cup C)$
@@
In Section 4 we saw that, even without Assumption 5.1, $\smile$$|^{n}$ has all the properties of an independence relation restricted to $M_{n}$ except, possibly, for the symmetry property. Now we have:
@@ 6. Connection to rosiness
    A formula $\varphi(\bar{x},\bar{a})$ (with all parameters listed by $\bar{a})$) strongly divides over $C$ if $\mathrm{tp}(\bar{a}/C)$ is nonalgebraic (i.e. has infinitely many realizations) and, for some $k<\omega$, the set of formulas $\{\varphi(\bar{x},\bar{a}^{\prime}):\mathrm{tp}(\bar{a}^{\prime}/C)=\mathrm{tp}(\bar{a}/C)\}$ is $k$-inconsistent (meaning that any set of $k$ of these formulas is inconsistent, in a model of $T^{\mathrm{eq}}$).
@@
Proof. The assumption that $\varphi(x,b,\bar{c})$ strongly divides over $C$ means that $\mathrm{tp}(b/C)$ is nonalgebraic and, for some $k<\omega$, $\{\varphi(x,b^{\prime},\bar{c}):\mathrm{tp}(b^{\prime}/C)=\mathrm{tp}(b/C)\}$ is $k$-inconsistent. As explained in [22, Remark 2.1.2], it follows from a compactness argument that there is $\theta(y,\bar{d})\in\mathrm{tp}(b/C)$ such that
@@
We say that $\smile$$|^{n}$ is symmetric restricted to $M_{n}$ if for all $a,b\in M_{n}$, and $C\subseteq M^{\mathrm{eq}}$, if $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}b$ then $b\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}a$.
@@
Proof. Let $m<\omega$, let $\bar{a}$ and $\bar{b}$ be finite sequences of elements from $M_{m}$, and let $C\subseteq M^{\mathrm{eq}}$. First we show that if $\mathrm{tp}(\bar{a}/\bar{b}C)$ thorn divides over $C$, then $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}\bar{b}$ for all $n\geq m$. For a contradiction, let $m\leq n<\omega$ and suppose that $\mathrm{tp}(\bar{a}/\bar{b}C)$ thorn divides over $C$ and $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}$. Then some $\varphi(\bar{x},\bar{b},\bar{c})\in\mathrm{tp}(\bar{a}/\bar{b}C)$, where $\bar{c}\in C$, thorn divides over $C$. This means that there is $D\supseteq C$ such that $\varphi(\bar{x},\bar{b},\bar{c})$ strongly divides over $D$.
The assumptions that $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}$ and that $\mathrm{rng}(\bar{a})\subseteq M_{n}$ together with the extension property of $\smile$$|^{n}$ (Lemma 4.8) implies that there is $\bar{a}^{\prime}$ such that $\mathrm{tp}(\bar{a}^{\prime}/\bar{b}C)=\mathrm{tp}(\bar{a}/\bar{b}C)$ and $\bar{a}^{\prime}\underset{\bar{b}C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$, so in particular $\mathrm{rng}(\bar{a}^{\prime})\subseteq M_{n}$. The assumption that $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}$ (and automorphism invariance of $\smile$$|^{n}$) gives $\bar{a}^{\prime}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}$. By transitivity of $\smile$$|^{n}$ we then get $\bar{a}^{\prime}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}D$. Due to the saturation assumption about $\mathcal{M}^{\mathrm{eq}}$ and since $\mathrm{tp}(\bar{a}/\bar{b}C)=\mathrm{tp}(\bar{a}^{\prime}/\bar{b}C)$ there is an elementary function $\sigma$, including $\bar{a}^{\prime}\bar{b}CD$ in its domain, such that $\sigma$ fixes $\bar{b}C$ pointwise and $\sigma(\bar{a}^{\prime})=\bar{a}$. By automorphism invariance of $\smile$$|^{n}$ we get $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}\sigma(D)$. As strong dividing is preserved by elementary functions, $\varphi(\bar{x},\bar{b},\bar{c})$ strongly divides over $\sigma(D)$. Without loss of generality, rename $\sigma(D)$ as $D$. Then $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}D$ and $\varphi(\bar{x},\bar{b},\bar{c})$ strongly divides over $D$.
From $\bar{a}^{\prime}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}D$ and monotonicity of $\smile$$|^{n}$ we get $\bar{a}\underset{D}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}D$ which means the same as $\bar{a}\underset{D}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}$.
Since $\mathcal{M}^{\mathrm{eq}}\models\varphi(\bar{a},\bar{b},\bar{c})$ and $\varphi(\bar{x},\bar{b},\bar{c})$ strongly divides over $D$ it follows from Lemma 6.2 that $\bar{b}\in\mathrm{acl}^{\mathrm{eq}}(\bar{a}D)\setminus\mathrm{acl}^{\mathrm{eq}}(D)$. Since $\bar{b}\in M_{n}$ it follows that $\bar{b}\in\mathrm{acl}^{n}(\bar{a}D)\setminus\mathrm{acl}^{n}(D)$. Since Assumption 6.5 and Proposition 5.14 imply that $\smile$$|^{n}$ has the symmetry property restricted to $M_{n}$, it follows from Lemma 6.4 that $\bar{a}\underset{D}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}\bar{b}$ and this contradicts the conclusion above that $\bar{a}\underset{D}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}\bar{b}$. Now we have proved that
(6.1) If $m\leq n$, $\bar{a},\bar{b}\in M_{m}$, and $\mathrm{tp}(\bar{a}/\bar{b}C)$ thorn divides over $C$, then $\bar{a}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}\bar{b}$.
@@
Proof. By Remark 2.8 it suffices to prove that there do not exist $a\in M^{\mathrm{eq}}$ and $b_{\alpha}\in M^{\mathrm{eq}}$, for $\alpha<\aleph_{1}$, such that $a\underset{(b_{i})_{i<\alpha}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\alpha}$ for all $\alpha<\aleph_{1}$. Towards a contradiction, suppose that there are $a\in M^{\mathrm{eq}}$ and $b_{\alpha}\in M^{\mathrm{eq}}$, for $\alpha<\aleph_{1}$, such that $a\underset{(b_{i})_{i<\alpha}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\alpha}$ for all $\alpha<\aleph_{1}$. Let $B=\{b_{\alpha}:\alpha<\aleph_{1}\}$. By Lemma 3.12, there is a countable $C\subseteq B$ such that $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ for all $n<\omega$. Let $\beta=\sup\{\alpha:b_{\alpha}\in C\}$. As $\aleph_{1}$ is a regular cardinal and $C$ is countable it follows that $\beta<\aleph_{1}$, so $\beta$ is a countable ordinal. Let $B_{\beta}=\{b_{\alpha}:\alpha\leq\beta\}$ so $B_{\beta}$ is countable and $C\subseteq B_{\beta}$. By the monotonicity of $\smile$$|^{n}$ we get $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ for all $n<\omega$. By assumption we have $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\beta+1}$, so Lemma 6.6 (which uses Assumption 6.5) implies that $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b_{\beta+1}$ for all sufficiently large $n$. Since $b_{\beta+1}\in B$ it follows that $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}B$ for all sufficiently large $n$. But this contradicts the earlier conclusion that $a\underset{B_{\beta}}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ for all $n<\omega$. $\square$
@@
Definition 6.13.
[7, 22] Let $p(\bar{x})$ be a complete type over some $A\subseteq M^{\mathrm{eq}}$ (and we assume that consistency is part of the definition of being a type).
  1. (1)
    We define the $U^{\text{\rm\th }}$-rank of $p(\bar{x})$ as follows:
    1. (a)
      $U^{\text{\rm\th }}(p(\bar{x})\geq 0$.
    2. (b)
      For every ordinal $\alpha$, $U^{\text{\rm\th }}(p(\bar{x}))\geq\alpha+1$ if there is $b\in M^{\mathrm{eq}}$ and a complete type $q(\bar{x})$ over $Ab$ such that $p(\bar{x})\subseteq q(\bar{x})$, $U^{\text{\rm\th }}(q(\bar{x}))\geq\alpha$, and $q(\bar{x})$ thorn forks over $A$.
    3. (c)
      For every limit ordinal $\beta$, $U^{\text{\rm\th }}(p(\bar{x}))\geq\beta$ if $U^{\text{\rm\th }}(p(\bar{x}))\geq\alpha$ for all ordinals $\alpha<\beta$.
    4. (d)
      For every ordinal $\alpha$, $U^{\text{\rm\th }}(p(\bar{x}))=\alpha$ if $U^{\text{\rm\th }}(p(\bar{x}))\geq\alpha$ and $U^{\text{\rm\th }}(p(\bar{x}))\not\geq\alpha+1$.
  2. (2)
    A theory is superrosy if for every complete type $p(x)$ over $\emptyset$ there is an ordinal $\alpha$ such that $U^{\text{\rm\th }}(p(x))=\alpha$. If in addition $U^{\text{\rm\th }}(p(x))<\omega$ for every complete type $p(x)$ over $\emptyset$, then we say that the theory has finite $U^{\text{\rm\th }}$-rank.
Since thorn forking has the extension property and since $U^{\text{\rm\th }}(\bar{a}/B)\leq U^{\text{\rm\th }}(\bar{a}/A)$ if $A\subseteq B$ (see e.g. [7, Theorem 2.7]) it follows from [7, Fact  4.4] that the definition of superrosy theory above is equivalent to the one given in [7, Section 4.1].
@@
If $T$ is $\omega$-categorical with geometric elimination of imaginaries and Assumption 6.5 holds for $n=0$, then $T$ is superrosy with finite $U^{\text{\rm\th }}$-rank.
Proof. Let $T$ be $\omega$-categorical with geometric elimination of imaginaries and suppose that Assumption 6.5 holds for $n=0$. It follows from Theorem 5.15 that $\smile$$|^{0}$ is an independence relation. Since $T$ has geometric elimination of imaginaries Lemma 4.9 implies that, for all $n<\omega$ and all $A,B,C\subseteq M^{\mathrm{eq}}$, $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}B$. Hence, for all $n<\omega$, $\smile$$|^{n}$ is an independence relation. It follows that Assumption 6.5 holds for all $n<\omega$, so Lemma 6.6 implies that if $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{\text{\rm\th }}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$ then $a\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$ for all sufficiently large $n$.